OpenStudy (anonymous):

Find the sum: (3x2 + 5x + 8) + (6x2 + 3x + 2)

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (whpalmer4):

Collect like terms. add up the x^2 terms, add up the x terms, add up the numeric terms. For example \[(x^2 + 2x + 1) + (3x^2 + 4x + 2) = (1+3)x^2 + (2+4)x + (1+2) = 4x^2 + 6x + 3\]

OpenStudy (anonymous):

Then what ?

OpenStudy (whpalmer4):

Then you're done! What do you get?

OpenStudy (anonymous):

3X²+2X+6 ?

OpenStudy (whpalmer4):

What is 3x^2 + 6x^2? What is 5x + 3x? What is 8 + 2? What do you get when you add those three answers together?

OpenStudy (anonymous):

\[(aX^{2}+bX+c)+(qX^{2}+rX+s)=(a+q)X^{2}+(b+r)X+(c+s)\]

OpenStudy (anonymous):

(3x2 + 5x + 8) + (6x2 + 3x + 2) 3x^2+ 6x^2= 9x^2 5x+3x=8x 8+2=10 So you get 9x^2 +8x +10